OpenStudy (anonymous):

√x^2negativex-12=x+3

OpenStudy (watchmath):

\(\sqrt{x^2-x-12}=x+3\) like that?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

When I tried putting the negative sign in front of the x and then minusing the 12 it thought I was trying to subtract the x

OpenStudy (watchmath):

\(\sqrt{(x+3)(x-4)}=(x+3)\) Square both sides \((x+3)(x-4)=(x+3)(x+3)\) \((x+3)((x-4)-(x+3))=0\) \((x+3)(-7)=0\) Hence \(x=-3\) is the only answer.

OpenStudy (anonymous):

Thank you !!!